OpenStudy (anonymous):

4x - (2x + 6) = 8x - 12 x = ?

alones (alones):

Hrmmm 4x−(2x+6)=8x−12 2x−6=8x−12 2x−6−8x=8x−12−8x −6x−6=−12 -6x−6+6=−12+6 −6x=−6 −6x/−6=−6/−6 x=1

rebeccaxhawaii (rebeccaxhawaii):

^^

alones (alones):

rebecca you forgot a nose ..like this ^.^

OpenStudy (anonymous):

5x - 8 = 2x + 1 with this one would i get rid of the 2x first? @AloneS

alones (alones):

First subtract 2x from both sides. 5x−8−2x=2x+1−2x 3x−8=1 that wouul be your firsdt step

OpenStudy (anonymous):

now i get rid of 8?

OpenStudy (anonymous):

the answer i got was x=3

alones (alones):

Yes that's hte answer! NIce Job<

alones (alones):

*<3 xP

OpenStudy (anonymous):

-3x = 2x + 15 do i subtract 2x from both sides again

OpenStudy (anonymous):

@AloneS

alones (alones):

Yes you do subtract first 2x from both sides. −3x−2x=2x+15−2x −5x=15 Step 2: Divide both sides by -5. −5x/−5=15/−5 x=−3

OpenStudy (anonymous):

=) thanks